OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. –x + 5y = 1 2x + 4y = –4 Answers: 2(–5y + 1) + 4y = –4 2x + 4(5y – 1) = –4 2x + 4(–5y + 1) = –4 2(5y – 1) + 4y = –4

OpenStudy (anonymous):

the last one i think

OpenStudy (anonymous):

because -x +5y = 1 is the same as -x = 1-5y x = 5y-1